OpenStudy (anonymous):

Solve y+3y=96 A. y= 4 B. 4y C. y=24 D. y=96

OpenStudy (anonymous):

@HWBUSTER00

OpenStudy (anonymous):

@RainbowBrony555

OpenStudy (anonymous):

I'm sorry. I'm bad at maths

OpenStudy (anonymous):

oh thats fine @RainbowBrony555

OpenStudy (anonymous):

when there is no number in front of y that means 1 so 1y+3y=96

OpenStudy (anonymous):

also @Rainbow_Dash_lov3 there is this helpful website i use when i cant figure stuff out for math http://www.cymath.com/answer.php?q=

OpenStudy (anonymous):

ok thanx @HWBUSTER00

OpenStudy (anonymous):

i think i figured out how to do this problem

OpenStudy (anonymous):

the answers provided were a. y=4 b. 4y c. y=24 d. y=96

OpenStudy (anonymous):

so i just chose a random one which was c. y=24

OpenStudy (anonymous):

and the question was y+3y=96

OpenStudy (anonymous):

yes thats correct because 4 / 96 is 24

OpenStudy (anonymous):

so i put 24 for the y

OpenStudy (anonymous):

and that would be 24+3*24=96

OpenStudy (anonymous):

so

OpenStudy (anonymous):

3*24=72+24=96 @HWBUSTER00

OpenStudy (anonymous):

right?